OpenStudy (anonymous):

Graph the following line by plotting the x and y intercepts: x + 3y = 6

OpenStudy (anonymous):

find x and y intercept x intercept: set y=0 x+3(0)=6 x+0=6 x=6 (6,0) y intercept: set x=0 (0)+3y=6 3y=6 divide each side by 3 y=2 (0,2)

OpenStudy (anonymous):